OpenStudy (anonymous):

Given points P(2,-1), Q(-4, 2), and M(5,3), find coordinates of a point K such that MK is perpendicular to PQ? Help Please !!

OpenStudy (anonymous):

I didn't compute it, but here's what you can do. Suppose (a,b) is the point K. Find the line that goes through the points M and K and its slope. If MK _|_ PQ, then their line slopes are inverse proportional and anti simetric.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What don't you understand?

OpenStudy (anonymous):

linear equation of PQ is (y-y1)/y2-y1)=(x-x1)/(x2-x1) so it's y=(-1/2)x if you wanna have a perpendicular line their slope must be negative and reverse m1*m2=-1 so the slope must be 2 if you have the slope and a point the linear equation is (y-y1)=m(x-x1) so it's (y-2)=2(x-5) so y=2x-7

OpenStudy (anonymous):

but there aren't any line that goes through M and be perpendicular of PQ here

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):